# Motivic Homotopy Groups of Spheres and Free Summands of Stably Free Modules

### v2

@@ Abstract.
Working over an algebraically closed field of characteristic 0, we show that the motivic stable homotopy groups of the sphere spectrum can be determined entirely from the motivic homotopy groups of the $p$-completed sphere spectra and the motivic cohomology of the ground field, except possibly for the 0 and 1-1-stems. Using this, we show that the complex realization maps from the motivic homotopy groups to the classical stable homotopy groups are an isomorphism in a range of bidegrees. We apply this to deduce that complex realization also induces isomorphisms on unstable homotopy groups for Stiefel varieties $V_{r}(\mathbb{A}^{n}_{k})$ in a range of bidegrees. This allows a complete solution of the question of when the projection map $V_{r}(\mathbb{A}^{n}_{k})\to V_{1}(\mathbb{A}^{n}_{k})$ admits a right inverse. Equivalently, this settles the question of when the universal stably-free module of type (n,n1)(n,n-1) admits a free summand of given rank.
@@
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2021-02603.
@@ 1. Introduction
In [9], it is proved that a stably free module $P$ of type (24m,24m1)(24m,24m-1) admits a free summand of rank $2$ if $R$ contains a field $k$ of characteristic 0, and of rank $3$ if that field has at most one quadratic extension. The method is to convert the problem to motivic homotopy theory: specifically, the injectivity of a realization map on the homotopy groups of Stiefel varieties. This problem is then solved by using recent excellent progress on the calculation of motivic homotopy sheaves of spheres—notably, [22] and [23]. These calculations of the homotopy sheaves are of a general nature, valid over any field.
@@
A great deal more is known about the motivic homotopy groups over an algebraically closed field of characteristic 0, using the motivic Adams and Adams–Novikov spectral sequences: [8, 13, 15, 26, 14] and elsewhere. The apparent obstacle to using these calculations in our work is that these sequences calculate the homotopy groups of $p$-completed objects, rather than of the sphere spectrum itself. In this paper, we explain how to surmount this obstacle using some algebra and the results of [2], and give a large range of bidegrees $(s,w)$ in which the motivic and classical homotopy groups coincide. Using this, and assuming our rings contain $\bar{\mathbb{Q}}$, we can then push the methods of [9] much further, fully solving a problem implicit in [20] of determining which universal examples of stably free modules admit rank-$r$ free summands.
1.1. Outline
In Section 2, we discuss the Ext-completion $\operatorname{Ext}(\mathbb{Q}/\mathbb{Z},A)$ of an abelian group $A$, which may be viewed as a completion at all primes. We also handle the variant where only a subset of the prime numbers is considered. We make no claims to originality about this section, but we need it as a reference for our arguments concerning the homotopy groups of $p$-completed motivic spectra.
Let $k=\bar{k}$ be an algebraically closed field of characteristic 0. The main body of the paper begins with Theorem 3.2, which says that in nonnegative weights $w$ and aside from the $s=0$-stem, the motivic stable homotopy group $\bm{\pi}_{s,w}(\mathbb{1})(k)$ of the sphere spectrum is isomorphic to the product over all primes $p$ of the groups $\bm{\pi}_{s,w}(\mathbb{1}^{\wedge}_{p})(k)$ of the $p$-completed spheres. A weaker version holds for negative weights: the completions may fail to detect the maximal divisible subgroup, which in most cases coincides with the motivic cohomology group $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$.
This implies that the extensive calculations of homotopy groups of the $p$-completed sphere amount to computations of the motivic homotopy groups $\bm{\pi}_{s,w}(\mathbb{1})(k)$ of the sphere spectrum when $s\neq 0$ and $w\geq 0$ (in fact, when $w\geq-1$).
An immediate consequence of Theorem 3.2 is that $\bm{\pi}_{s,w}(\mathbb{1})(k)$ is isomorphic to the classical stable homotopy group of the sphere spectrum, $\pi_{s}(\mathbb{S})$, when $s\neq 0$ and $0\leq w\leq\frac{1}{2}(s+1)$—this relies on isomorphisms of the $p$-completions, which are known from [10] and [25].
@@ 1.1. Outline
There are two ingredients in the proof of Theorem 3.2. First we consider the algebraic properties of the $\operatorname{Ext}$-completions $\operatorname{Ext}(\mathbb{Z}/(p^{\infty}),\bm{\pi}_{s,w}(\mathbb{1})(k))$, which agrees with $\bm{\pi}_{s,w}(\mathbb{1}^{\wedge}_{p})(k)$ except when $s=0$. Finiteness of $\prod_{p}\bm{\pi}_{s,w}(\mathbb{1})(k)$ is now sufficient to prove that the completion map belongs in a split short exact sequence
$0\to D_{s,w}\to\bm{\pi}_{s,w}(\mathbb{1})(k)\to\prod_{p}\bm{\pi}_{s,w}(\mathbb{1}^{\wedge}_{p})(k)\to 0$
@@
where $D_{s,w}$ is the subgroup of divisible elements in $\pi_{s,w}(\mathbb{1})(k)$. Except when $s\in\{0,-1\}$, this kernel is observed to be a $\mathbb{Q}$-vector space. We then rely on [2], which draws on [7], for the identification
$\mathbb{Q}\otimes_{\mathbb{Z}}\bm{\pi}_{s,w}(\mathbb{1})(k)\cong\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$
@@
Having proved that the motivic and classical stable homotopy groups of spheres agree, we then exploit the $\mathbb{P}^{1}$-Freudenthal theorem of [3] and an induction argument to prove that complex realization induces isomorphisms on homotopy groups of Stiefel varieties. Here we suppose $k$ is embedded as a subfield of $\mathbb{C}$. If $V_{r}(\mathbb{A}^{n})$ denotes $\operatorname{GL}(n)/\operatorname{GL}(n-r)$ and $W_{r}(\mathbb{C}^{n})$ denotes its complex analogue U(n)/U(nr)U(n)/U(n-r), then the realization map
(1) $\bm{\pi}_{d+e,e}(V_{r}(\mathbb{A}^{n}))(k)\to\pi_{d+e}(W_{r}(\mathbb{C}^{n}))$
is an isomorphism for a set of values of $(d,e)$ depending on $(n,r)$, and an injection for a larger set. The precise statements are Theorems 4.4 and 4.5.
These theorems combine with [9] to imply that right inverses for projection maps $V_{r}(\mathbb{A}_{\bar{\mathbb{Q}}}^{n})\to V_{1}(\mathbb{A}_{\bar{\mathbb{Q}}}^{n})$ exist if and only if right inverses exist for the corresponding projections $W_{r}(\mathbb{C}^{n})\to W_{1}(\mathbb{C}^{n})$. Following [1], it is known that the latter exists if and only if $b_{r}\mid n$ where $b_{r}$ is the $r$th James number. We deduce Theorem 5.1, saying the same thing for the Stiefel variety.
Finally, we establish Corollary 5.2, which says that if $R$ is a commutative ring containing an algebraic closure of $\mathbb{Q}$, and if $P$ is a stably free $R$-module satisfying $P\oplus R\cong R^{n}$, and if $b_{r}\mid n$, then $P$ admits a decomposition $R^{r}\oplus Q\cong P$. This settles a question left open by [20, Thm. 6.5] in the case where the ground field contains an algebraic closure of $\mathbb{Q}$.
1.2. Remarks and conventions
We use the symbol $\mathbb{1}$ for the motivic sphere spectrum, and $\mathbb{S}$ for the classical sphere spectrum. This usage is borrowed from [23], and is justified by convenience.
When we work with motivic spaces or spectra over a base field $k$, we use a boldface $\bm{\pi}_{\ast}(X)$ for the homotopy sheaves. If $X$ is a motivic object, then $\pi_{\ast}(X)=\bm{\pi}_{\ast}(X)(k)$ is used for the group (or pointed set) of global sections.
@@ 1.2. Remarks and conventions
which we will call the “stem-weight” bigrading, and the convention of [12]:
@@
which we will call the “coweight-weight” bigrading. Stem-weight bigrading is particularly well suited to motivic Adams and Adams–Novikov spectral sequence calculations over $\mathbb{C}$, and is what is used in the extensive literature on this topic. On the other hand, coweight-weight bigrading is better for use in problems concerning motivic connectivity, for instance, the motivic Freudenthal suspension theorem of [3]. As a consequence, we use both conventions in this paper. In practice, this means that the notation $c+w\alpha$ should be read as $(c+w,w)$. Here is the same homotopy group written in two ways:

### v3

@@ Abstract.
Working over an algebraically closed field $k$ of characteristic 0, we show that the motivic stable homotopy groups of the sphere spectrum can be determined entirely from the motivic homotopy groups of the $p$-completed sphere spectra and the motivic cohomology of the ground field, except possibly for the 0 and 1-1-stems. Using this, we show that the complex realization maps from the motivic homotopy groups to the classical stable homotopy groups are isomorphisms in a range of bidegrees. We apply this to deduce that complex realization also induces isomorphisms on unstable homotopy groups for Stiefel varieties $V_{r}(\mathbb{A}^{n}_{k})$ in a range of bidegrees. We use this to determine when the projection map $V_{r}(\mathbb{A}^{n}_{k})\to V_{1}(\mathbb{A}^{n}_{k})$ admits a right inverse, settling the question of when the universal stably-free module of type (n,n1)(n,n-1) admits a free summand of given rank.
@@
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2021-02603. We acknowledge that this research was supported in part by the Pacific Institute for the Mathematical Sciences (PIMS) under CRG 41.
@@ 1. Introduction
In [undefi], we proved that a stably free module $P$ of type (24m,24m1)(24m,24m-1) admits a free summand of rank $2$ if $R$ contains a field $k$ of characteristic 0, and of rank $3$ if that field has at most one quadratic extension. The method is to convert the problem to motivic homotopy theory: specifically, whether a realization map relating the motivic and classical homotopy groups of Stiefel varieties is injective. This problem is then solved by using recent excellent progress on the calculation of the motivic homotopy sheaves of spheres—notably, [undefae] and [undefaf]. Their calculations are of a general nature, valid over any field, but limited in the the bidegrees to which they apply.
@@
Let $k=\bar{k}$ be an algebraically closed field of characteristic 0. A great deal more is known about the motivic homotopy groups over $k$ than in general, using the motivic Adams and Adams–Novikov spectral sequences: [undefg, undefp, undefq, undefr, undefai] and other work. The apparent obstacle to using these calculations in our work is that these spectral sequences calculate the homotopy groups of $p$-completed objects, rather than of the sphere spectrum itself. We surmount this obstacle and consequently generalize the results of [undefi] to all $r$ over the base field $k$.
We write $\mathbb{1}$ for the motivic sphere spectrum, and $\bm{\pi}_{s,w}(\mathbb{1})$ for the bigraded motivic homotopy sheaf in bigrading $(s,w)$. We use a regular font to denote the group of global sections of this sheaf, so that
$\pi_{s,w}(\mathbb{1})=\bm{\pi}_{s,w}(\mathbb{1})(k).$
We will also write $\mathsf{M}\mathbb{Z}$ for the motivic Eilenberg–MacLane spectrum. This shown to be a commutative ring spectrum in [undefh, Example 3.4]. In particular there is a unit map $\mathbb{1}\to\mathsf{M}\mathbb{Z}$ inducing maps $\pi_{s,w}(\mathbb{1})\to\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$ on homotopy groups, which we also call “unit maps”. They may also be considered Hurewicz maps if we use $\mathsf{M}\mathbb{Z}$ to represent a homology theory: $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))=\mathrm{H}_{s}(\operatorname{Spec}k;\mathbb{Z}(w))$.
The paper establishes four main theorems. Our first concerns the relationship between the sphere spectrum and its $p$-completions. We write $\mathcal{P}$ for the set of all prime numbers.
Theorem 1.1.
Let $s,w$ be integers satisfying $s\neq 0$. Then the comparison map
is split surjective, and the kernel is the subgroup of divisible elements. Additionally, if $s\neq-1$, then the kernel is isomorphic to the motivic cohomology group $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$ and the inclusion of the kernel is split by the unit map $\pi_{s,w}(\mathbb{1})\to\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$.
This theorem assures us that the Adams and Adams–Novikov-spectral sequences do, with minor exceptions, calculate the motivic homotopy groups (i.e., the global sections of the homotopy sheaves) of $\mathbb{1}$ over $k$. The kernel $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$ is either known or conjectured to vanish in a large range of bidegrees.
Notation 1.2.
We write $\mathbb{S}$ for the classical sphere spectrum, to distinguish it from the motivic $\mathbb{1}$. This notation is adopted from [undefaf], and is justified by its usefulness.
Among the implications of Theorem 1.1 is that complex realization induces an isomorphism
(1) $\pi_{s,w}(\mathbb{1})\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{s}(\mathbb{S}),\text{\quad when $-1\leq w\leq\frac{1}{2}s+1$ and $1\leq s$},$
where the target is the ordinary stable homotopy group of the classical sphere spectrum. This is because the maps $\pi_{s,w}(\mathbb{1}^{\wedge}_{p})\to\pi_{s}(\mathbb{S}^{\wedge}_{p})$ are isomorphisms in this range; see for instance [undefj] when $p=2$ and [undefah] when $p>2$. We give a fuller account of what we now know about $\pi_{s,w}(\mathbb{1})$ in Subsection 3.1 below.
@@
There are two ingredients in the proof of Theorem 1.1. First we consider the algebraic properties of the $\operatorname{Ext}$-completions $\operatorname{Ext}(\mathbb{Z}/(p^{\infty}),\pi_{s,w}(\mathbb{1}))$, which agree with $\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$ except when $s=0$. Finiteness of $\prod_{p}\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$ is now sufficient to prove that the completion map belongs in a split short exact sequence
$0\to D_{s,w}\to\pi_{s,w}(\mathbb{1})\to\prod_{p}\pi_{s,w}(\mathbb{1}^{\wedge}_{p})\to 0$
@@
where $D_{s,w}$ is the subgroup of divisible elements in $\pi_{s,w}(\mathbb{1})(k)$. Except when $s\in\{0,-1\}$, this kernel is observed to be a $\mathbb{Q}$-vector space. We then rely on [undefa], which draws on [undefe], for the identification
$\mathbb{Q}\otimes_{\mathbb{Z}}\pi_{s,w}(\mathbb{1})\cong\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$
@@
Remark 1.3.
The arguments we use here are prefigured by the proof of [undefm, Lemma 2.10]. Although that lemma is concerned with lifting one specific element $\mu_{9}\in\pi_{9,5}(\mathbb{1}^{\wedge}_{2})$ to $\pi_{9,5}(\mathbb{1})$, the proof that is given there is considerably more general and agrees in outline with much of our argument for Theorem 1.1. We are grateful to Jens Hornbostel for pointing this out to us.
Since our ambition is to establish geometric and algebraic results, we pass from stable to unstable homotopy theory. Our tool here is the motivic Freudenthal suspension theorem of [undefb]. The unstable analogs of the isomorphisms (1) are recorded in Proposition 4.3.
For application to projective modules, the most important motivic spheres are the spheres $S^{2n-1,n}\simeq\operatorname{GL}(n)/\operatorname{GL}(n-1)$, which form part of the larger family of Stiefel varieties $V_{r}(\mathbb{A}^{n}_{k})=\operatorname{GL}(n)/\operatorname{GL}(n-k)$. Here $\operatorname{GL}(n-k)$ is embedded in $\operatorname{GL}(n)$ by block summation.
(2) $A\mapsto\begin{bmatrix}I_{k}&0\\ 0&A\end{bmatrix}$,
where $I_{k}$ is the $k\times k$ identity matrix. The Stiefel varieties have the property that they are the base spaces for universal stably free modules, as described in [undefac] and in [undefi, Sec. 3]. Specifically, a morphism of schemes
$\phi:\operatorname{Spec}R\to V_{r}(\mathbb{A}_{k}^{n})$
corresponds to an $R$-module $P$ equipped with an isomorphism $P\oplus R^{r}\to R^{n}$.
Let $U(r)$ denote the unitary group of $r\times r$ matrices. The complex Stiefel manifolds are the quotient manifolds $W_{r}(\mathbb{C}^{n})=U(n)/U(n-k)$, where the embeddings follow the same pattern as in (2).
Our second and third main theorems constitute a pair, and they deal with the calculation of the unstable homotopy groups of the Stiefel varieties. Before we state these theorems, we discuss some notation.
Notation 1.4.
@@
which we will call the “stem-weight” bigrading, and the convention of [undefo]:
@@
which we will call the “coweight-weight” bigrading. Stem-weight bigrading is particularly well suited to motivic Adams and Adams–Novikov spectral sequence calculations over $\mathbb{C}$, and is what is used in the extensive literature on this topic. On the other hand, coweight-weight bigrading is better for use in problems concerning motivic connectivity, for instance, the motivic Freudenthal suspension theorem of [undefb]. As a consequence, we use both conventions in this paper. In practice, this means that the notation $c+w\alpha$ should be read as $(c+w,w)$. Here is the same homotopy group written in both conventions:
@@
Complex realization induces maps of homotopy groups
$\pi_{d+e\alpha}(V_{r}(\mathbb{A}^{n}))=\pi_{d+e,e}(V_{r}(\mathbb{A}^{n}))\to\pi_{d+e}(W_{r}(\mathbb{C}^{n})).$
One may inquire how close these maps are to being isomorphisms. The complexity of our answer means that it is clearer to divide it between two theorems. The first establishes surjectivity of the map in a range of bidegrees, which specializes to establish an isomorphism in some cases.
Theorem 1.5.
  1. (1)
    $r\leq n-2$;
  2. (2)
    $d\leq 2n-2r-3$;
  3. (3)
    $e\leq d+4-r$;
  4. (4)
    $2n\leq e+d$;
The complex realization map
$\pi_{d+e\alpha}(V_{r}(\mathbb{A}^{n}))\to\pi_{d+e}(W_{r}(\mathbb{C}^{n}))$
If additionally $n-1\leq e$, then the realization map is an isomorphism.
This has the disadvantage that the admissible $e$-values are starkly constrained when $r$ is large by the inequality $e\leq d+4-r$. The second theorem establishes injectivity without this constraint.
Theorem 1.6.
  1. (1)
    $r\leq n-2$;
  2. (2)
    $d\leq 2n-2r-3$;
  3. (3)
    $e\leq d+3$;
  4. (4)
    $2n\leq e+d$;
  5. (5)
    $n-1\leq e$.
The realization map
$\pi_{d+e\alpha}(V_{r}(\mathbb{A}^{n}))\to\pi_{d+e}(W_{r}(\mathbb{C}^{n}))$
is injective.
The methods of proofs of both theorems are the same, induction on $r$ using homological algebra in the mode of the five lemma, the base case of the spheres being supplied by Proposition 4.3.
Our fourth main theorem concerns the existence of a right inverse (or section) to a projection map $\rho:V_{r}(\mathbb{A}^{n}_{k})\to V_{1}(\mathbb{A}^{1}_{k})$. One construction of $\rho$ is as follows: assuming $1\leq r\leq n$ are integers, there are canonical embedding homomorphisms $\operatorname{GL}(n-r)\to\operatorname{GL}(n-1)\to\operatorname{GL}(n)$. Consequently there is a canonical map of quotients
$\operatorname{GL}(n)/\operatorname{GL}(n-r)=V_{r}(\mathbb{A}^{n}_{k})\overset{\rho}{\to}V_{1}(\mathbb{A}^{1}_{k})=\operatorname{GL}(n)/\operatorname{GL}(n-1).$
If one views $V_{r}(\mathbb{A}^{n}_{k})$ as representing stably free modules, the map $\rho$ corresponds to the construction that takes a module $P$ satisfying $P\oplus R^{r}\cong R^{n}$ and produces a module $Q=P\oplus R^{r-1}$, which satisfies $Q\oplus R\cong R^{n}$. In this guise, the map $\rho$ figures heavily in [undefac], where it is named “PP”, a letter we use for a projective module. In [undefac, Thm. 6.5], the existence of a right inverse for $\rho$ is ruled out except possibly when $n$ is divisible by a certain positive integer $b_{r}$, the $r$-th James (or Atiyah–Todd) number. Our last main theorem is the converse statement over $\bar{\mathbb{Q}}$. We state it as a biconditional, but only the “if” direction is new.
Theorem 1.7.
Let $n$ be a positive integer and suppose $2\leq r\leq n-2$. Then
has a right inverse if and only if $b_{r}\mid n$.
The restriction from $k$, assumed algebraically closed of characteristic 0, to $\bar{\mathbb{Q}}$ is no restriction at all; a right inverse constructed over $\bar{\mathbb{Q}}$ induces a right inverse over any field $k$ containing an algebraic closure of $\mathbb{Q}$.
The proof relies on work done in [undefi, §4 and §7], which we do not repeat in full here. The existence of a right inverse is equivalent to an ostensibly weaker condition: the induced map of homotopy groups
$\rho_{*}:\pi_{n-1+n\alpha}(V_{r}(\mathbb{A}^{n}_{\bar{\mathbb{Q}}}))\to\pi_{n-1+n\alpha}(V_{1}(\mathbb{A}^{n}_{\bar{\mathbb{Q}}}))$
is surjective. The equivalence is a statement that a morphism in the homotopy category can be lifted to a map of schemes: the variety $V_{1}(\mathbb{A}^{n}_{\bar{\mathbb{Q}}})$ is homotopy equivalent to $S^{n-1+n\alpha}$, so that $\pi_{n-1+n\alpha}(X)$ is the set of maps $V_{n-1+n\alpha}(\mathbb{A}^{n}_{\bar{\mathbb{Q}}})\to X$ in the motivic homotopy category.
In particular, any obstruction to the existence of a right inverse is witnessed by a nonzero cokernel for $\rho_{*}$, consisting of some elements of $\pi_{n-2+n\alpha}(V_{r-1}(\mathbb{A}^{n-1}_{\bar{\mathbb{Q}}}))$. This group is in the range of bidegrees covered by Theorem 1.6, so that any obstruction can be detected in the classical group $\pi_{2n-2}(W_{r-1}(\mathbb{C}^{n-1}))$. This allows the reduction of the whole problem to its classical analogue, the question of whether a right inverse exists for the projection $\rho:W_{r}(\mathbb{C}^{n})\to W_{1}(\mathbb{C}^{n})$. This was settled by Adams and Walker in [undef]: the answer being that a right inverse exists whenever $b_{r}\mid n$.
The new implication in Theorem 1.7 admits a translation into the language of stably free modules.
Corollary 1.8.
$P\oplus R\cong R^{n}$
$P\cong Q\oplus R^{r-1}.$
This corollary follows directly from Theorem 1.7 using [undefi, Prop. 4.4].
We sketch the argument used in [undefi] for the benefit of the reader who would prefer to keep reading this paper rather than stopping to refer to that one. Since $b_{r}\mid n$, there is a right-inverse map $\psi:V_{1}(\mathbb{A}^{n}_{\bar{\mathbb{Q}}})\to V_{r}(\mathbb{A}^{n}_{\bar{\mathbb{Q}}})$. We know that $V_{r}(\mathbb{A}^{n}_{\bar{Q}})$ represents $R$-modules $Q$ equipped with an isomorphism $Q\oplus R^{r}\cong R^{n}$, so that the map $\psi$ allows one to take any $P$ satisfying $P\oplus R\cong R^{n}$ and to produce a corresponding $Q$. The relation $\rho\circ\psi=\mathrm{id}$ implies that $P\cong Q\oplus R^{r-1}$.
1.1. Outline
In Section 2, we discuss the Ext-completion $\operatorname{Ext}(\mathbb{Q}/\mathbb{Z},A)$ of an abelian group $A$. This may be viewed as a completion at all primes. We also handle the variant where only a subset of the prime numbers is considered. We make no claims to originality about this section, but we need it as a reference for our arguments concerning the homotopy groups of $p$-completed motivic spectra.
Section 3 is devoted to applying these technical results to the sphere spectrum, proving Theorem 1.1 and collecting some of its immediate consequences, which include Corollary 3.3, which tells us that complex realization induces an isomorphism of homotopy groups in a range of bidegrees. Subsection 3.1, which contains no new results, is an incomplete guide to the groups $\pi_{s,w}(\mathbb{1})$ over an algebraically closed field of characteristic 0, included in the hope that the reader may find it useful.
Having proved that the motivic and classical stable homotopy groups of spheres agree in some bidegrees, in Section 4 we exploit the $\mathbb{P}^{1}$-Freudenthal theorem of [undefb] and an induction argument to prove that complex realization induces isomorphisms on unstable homotopy groups of motivic spheres and of Stiefel varieties in certain ranges. In Figure 2, we give a visual presentation of some of what we know about the unstable homotopy groups of the spheres.
In Section 5, we combine what we have proved so far with [undefi] to deduce that right inverses for projection maps $V_{r}(\mathbb{A}_{\bar{\mathbb{Q}}}^{n})\to V_{1}(\mathbb{A}_{\bar{\mathbb{Q}}}^{n})$ exist if and only if right inverses exist for the corresponding projections $W_{r}(\mathbb{C}^{n})\to W_{1}(\mathbb{C}^{n})$, thereby establishing our last main theorem, Theorem 1.7.